OpenStudy (anonymous):

Help me solve this? I'll give a medal and fan, ha. tan(3x)(tan x -1) = 0 interval [0,2p)

OpenStudy (kirbykirby):

Since you have 2 factors, then you have that \(\tan(3x)=0\) or \(\tan x-1=0 \implies \tan x=1\) This implies: \(\arctan(0)=3x\) or \(\arctan(1)=x\) Now, \(\arctan(0)=0\) and \(\arctan(1)=\dfrac{\pi}{4}\) But the tangent function is periodic over \(n \pi, ~~n\in \mathbb{Z}\) So you should actually have: \(3x=0+n\pi\) \(x=\dfrac{\pi}{4}+n\pi\) Find all of these values for n=0,1,2,.. etc. such that you stay within the given interval \([0, 2\pi)\)

OpenStudy (anonymous):

Thank you.

OpenStudy (kirbykirby):

:)